OpenStudy (anonymous):

Assuming 2^10 is approximately 1000, which one of the following best approximates 2^40? A. 4000 B. 10,000,000 C. 10^12 D. 1000^30

OpenStudy (asnaseer):

think of how to represent \(2^{40}\) in terms of \(2^{10}\)

OpenStudy (asnaseer):

using this rule:\[(x^a)^b=x^{ab}\]

OpenStudy (anonymous):

1000^4?

OpenStudy (asnaseer):

thats it!

OpenStudy (anonymous):

Which of my multiple choice answers would match that though?

OpenStudy (asnaseer):

now do the same with 1000. write it as \(10^a\) - what should 'a' be equal to here?

OpenStudy (anonymous):

10^3

OpenStudy (asnaseer):

perfect! so now you have:\[1000^4=(10^3)^4=???\]

OpenStudy (asnaseer):

use the same rule

OpenStudy (anonymous):

Ah. My answer is C. 10^12. That was easy. Thank you so much!

OpenStudy (asnaseer):

great! and yw :)